# Probing anomalous Higgs boson couplings in Higgs plus jet production at NLO QCD with full $m_{t}$-dependence

### v1

@@ 1 Introduction
Run 2 of the LHC was very successful in establishing the Higgs boson couplings to vector bosons and heavy fermions. With Run 3 and the high-luminosity upgrade the precision of the coupling measurements will further increase. Therefore it is important to have good control over the theoretical uncertainties, in particular to distinguish BSM effects from effects due to insufficient higher order corrections.
@@
It has been noticed some time ago Harlander:2013oja ; Azatov:2013xha ; Banfi:2013yoa ; Grojean:2013nya ; Schlaffer:2014osa ; Dawson:2014ora ; Buschmann:2014sia ; Langenegger:2015lra ; Azatov:2016xik ; Lindert:2018iug ; DiMicco:2019ngk that the Higgs boson $p_{T}$-spectrum is an important observable to constrain both the Yukawa-couplings to top (and bottom) quarks as well as effective gluon-Higgs couplings due to interactions with unknown particles at higher scales, leading to operators that are not present in the Standard Model (SM). The impact of such effective operators on the $p_{T,H}$ spectrum in Higgs+jet production, including NLO QCD corrections in the heavy top limit (HTL), has been investigated in Refs. Maltoni:2016yxb ; Grazzini:2016paz ; Grazzini:2018eyk ; Battaglia:2021nys ; Maltoni:2024dpn , where the analysis of Refs. Battaglia:2021nys ; Maltoni:2024dpn also includes investigations of effects due to the renormalisation group running of the Wilson coefficients.
@@
In the SM, the LO one-loop amplitude has been calculated in Ref. Baur:1989cm , NLO results beyond the HTL were first obtained for the high transverse momentum region Melnikov:2016qoc ; Lindert:2018iug . The NLO result for general kinematics has first been obtained numerically Jones:2018hbb ; Becker:2020rjp ; Neumann:2018bsx , the top quark mass effects have been studied in detail in Ref. Chen:2021azt .
@@
On the analytic side, after the relevant two-loop master integrals have become available Bonciani:2016qxi ; Bonciani:2019jyb ; Frellesvig:2019byn , the full NLO calculation based on analytic results for the two-loop integrals has been completed in Ref. Bonciani:2022jmb , including both top and bottom masses as well as a comparison of the on-shell and $\overline{\mathrm{MS}}$ schemes to renormalise the top quark mass. Two-loop bottom quark mass effects also have been studied in Ref. Pietrulewicz:2023dxt . Mixed QCD-EW corrections have been presented in Refs Bonetti:2020hqh ; Becchetti:2021axs ; Bonetti:2022lrk , partial EW corrections are considered in Refs. Becchetti:2018xsk ; Davies:2023npk ; Haisch:2024nzv . NNLO results for Higgs plus jet production in the HTL are available already since some time ago Boughezal:2015dra ; Boughezal:2015aha ; Caola:2015wna ; Chen:2016zka ; Campbell:2019gmd ; Chen:2019wxf ; Chen:2021ibm . Higgs production in association with one or multiple high-energy jets is available within the HEJ framework Andersen:2023kuj ; Andersen:2022zte . In Ref. Liu:2024tkc , light quark mediated Higgs boson production in association with a jet at NNLO and beyond is considered in the framework of resummation.
@@
The Higgs boson transverse momentum spectrum for boosted Higgs bosons already has been used by the experimental collaborations to place limits on anomalous top-Higgs and gluon-Higgs couplings ATLAS:2019lwq ; ATLAS:2021tbi ; CMS:2020zge ; CMS:2024jbe . The latter can be parameterised by $c_{t}$ and $c_{g}$ in Higgs Effective Field Theory (HEFT), also called Electroweak Chiral Lagrangian Feruglio:1992wf ; Burgess:1999ha ; Grinstein:2007iv ; Contino:2010mh ; Alonso:2012px ; Buchalla:2013rka . These couplings also enter inclusive Higgs boson production in gluon fusion, which is known to agree with the SM prediction to a level approaching 5%. Therefore, these anomalous couplings are fairly well constrained already (also from other processes such as $t\mathrm{bar}{t}H$ production for the case of the Higgs-top coupling Hartland:2019bjb ; Celada:2024mcf ). However, it is well known that there is a degeneracy between $c_{t}$ and $c_{g}$ when considering only inclusive Higgs production, which is lifted when considering the $p_{T}$-spectrum of the Higgs boson at large transverse momenta Grojean:2013nya ; Schlaffer:2014osa ; Grazzini:2016paz ; Grazzini:2018eyk . Up to now, the effects of these operators have not yet been studied in combination with NLO corrections to Higgs+jet production including the full top quark mass dependence. However, both the SM top quark mass effects as well as these anomalous couplings affect the tail of the $p_{T,H}$-distribution considerably. Therefore it is important to study in detail the interplay of both, higher order QCD corrections and potential effects of new physics in an EFT framework.
@@
In this work we would like to address this point and investigate the effects of these anomalous couplings on the Higgs+jet cross section and transverse momentum distributions, based on a calculation of the full NLO QCD corrections in the SM Jones:2018hbb . Working in HEFT rather than Standard Model Effective Field Theory (SMEFT) Buchmuller:1985jz ; Grzadkowski:2010es ; Brivio:2017vri ; Isidori:2023pyp , the leading effects of new physics residing at higher energy scales impacting this process can be parameterised by $c_{t}$ and $c_{g}$, while the chromomagnetic operator, pertaining to the class of loop-generated operators Arzt:1994gp ; Brivio:2017vri ; Buchalla:2022vjp ; Isidori:2023pyp , is considered as subleading and therefore will be considered in subsequent work, together with other subleading operators.
@@
The structure of this paper is as follows: In Section 2 we describe how the anomalous couplings relate to inclusive Higgs production and how they affect the large-$p_{T}$ spectrum of the Higgs boson. We also give some technical details about the calculation. Section 3 is dedicated to the description of phenomenological results, providing heat maps that show the effects of the anomalous couplings on the total cross section and discussing the effect of some HEFT benchmark points on the Higgs boson transverse momentum spectrum, before we conclude in Section 4.
@@ 2.1 Framework of the calculation
In the SM, $c_{t}=1$ and $c_{g}=0$. We assume that the anomalous couplings are induced by new physics interactions at a scale $\Lambda$ considerably larger than the electroweak scale. While the process $pp\to H+$jet is loop induced in the SM, the second part of the Lagrangian now also introduces effective tree level interactions. The factor proportional to $\alpha_{s}/8\pi$ indicates that these interactions are stemming from loops of heavy particles that have been integrated out to arrive at the effective Higgs-gluon coupling. In the region $2m_{t}\lesssim\sqrt{\hat{s}}\lesssim\Lambda$ the top quark loops are resolved while the heavier particles in the loop generate the effective point-like Higgs-gluon interaction. The coefficient $c_{t}$ is a modification factor of the top Yukawa coupling, which can arise for example by mixing with heavy top partners Grojean:2013nya ; Banfi:2019xai .
The chromomagnetic operator $\sim\mathrm{bar}Q_{L}\sigma^{\mu\nu}T^{a}G_{\mu\nu}^{a}\tilde{\phi}t_{R}$ can only be generated through contracted loops Arzt:1994gp ; Buchalla:2022vjp ; Isidori:2023pyp in weakly coupled, renormalisable UV completions. Sticking to such extensions of the SM, it comes with a loop suppression factor $\sim 1/(16\pi^{2})$ and therefore we do not include it here. Furthermore, we do not include any 4-quark operators nor CP-violating operators.
@@
where $\mathcal{M}_{f}$ denotes the parts of the amplitude where the Higgs boson couples to a top quark, and $\mathcal{M}_{Hg}$ the amplitude parts containing an effective point-like Higgs-gluon interaction. Note that the NLO amplitude can also contain both, a top quark loop and a point-like Higgs-gluon interaction. However, in these diagrams the top quarks couple only to gluons, with an SM coupling. We use $\kappa_{g}=\frac{3}{2}c_{g}$ such that the heavy top limit corresponds to $\kappa_{g}\to 1$ and $c_{t}=0$. The total cross section for $pp\to H+\rm{jet}$ can be written as a quadratic polynomial in $c_{t}$ and $c_{g}\,$, both at LO and at NLO.
Integrating over the jet momenta, the total inclusive cross section for Higgs boson production in gluon fusion should be retrieved. As is well known, due to the “Higgs Low Energy Theorem” Ellis:1975ap ; Vainshtein:1980ea ; Dawson:1989yh , the total cross section for Higgs production in gluon fusion is rather insensitive to the masses of heavy particles circulating in the loop. This is also reflected in the fact that, at energy scales below $2m_{t}$, the inclusive Higgs production cross section is approximated very well by the heavy top limit. An extra high-$p_{T}$ jet can serve as a handle to resolve heavy quark loops, therefore new physics effects could show up in the tail of the Higgs $p_{T}$-distribution.
@@
As the measured total Higgs production cross section in gluon fusion agrees very well with the SM result, the relation $(c_{t}+\kappa_{g})^{2}=1$ should be fulfilled to about 10% level, assuming that subleading operators do not have a drastic effect, which would lead to more freedom in the relation between $c_{t}$ and $\kappa_{g}$. In the pure heavy top limit, the proportionality of the cross section to $(c_{t}+\kappa_{g})^{2}$ is fulfilled exactly, also for the NLO amplitudes, because there are no diagrams at NLO which contain both $c_{t}$ and $c_{g}$ simultaneously, such that the heavy top limit of the full SM NLO amplitudes gives exactly the amplitudes proportional to $\kappa_{g}$. This degeneracy is broken in the Higgs boson transverse momentum spectrum because as $p_{T,H}$ increases, the top quark loops start to become resolved and therefore the kinematic behaviour of the contribution proportional to $c_{t}$ is different from the one in the HTL for large values of $p_{T,H}$. On the other hand, as the differential cross section decreases rapidly with $p_{T,H}$, the effects of anomalous couplings on the total cross section should be small as long as the relation $(c_{t}+\kappa_{g})^{2}=1$ is fulfilled. Therefore it is useful to consider the cross section for $pp\to H+\rm{jet}$ as a function of the cut on $p_{T,H}$ Grojean:2013nya ; Schlaffer:2014osa :
@@
where the coefficients $\delta_{i}$ depend on the cut $p_{T,H}^{\mathrm{min}}$. For small $p_{T,H}^{\mathrm{min}}$ the coefficients $\delta_{i}$ at LO are very small, modifying the cross section in the permille to percent range below $p_{T,H}\sim 350~{}\mathrm{G}\mathrm{eV}$ Grojean:2013nya . However, recent LHC measurements have reached transverse momentum regions beyond $600~{}\mathrm{G}\mathrm{eV}$ CMS:2024jbe ; CMS:2020zge ; ATLAS:2019lwq ; ATLAS:2023jdk . Furthermore, the study of Refs. Grojean:2013nya ; Schlaffer:2014osa was at LO only, and the one of Refs. Grazzini:2016paz ; Maltoni:2016yxb ; Grazzini:2018eyk is based on the heavy top limit when going beyond LO. In Section 3 we will investigate how the anomalous couplings modify the large-$p_{T,H}$ spectrum at NLO with full top quark mass dependence.
@@ 2.2 Technical details
The leading order amplitudes in the full theory as well as the amplitudes involving $c_{g}$ were implemented analytically, relying on Ref. Baur:1989cm , while the one-loop real radiation contribution and the two-loop virtual amplitudes rely on semi-numerical evaluations. As a cross-check we also generated the Born amplitudes with GoSam Cullen:2011ac ; GoSam:2014iqq using the UFO Degrande:2011ua ; Darme:2023jdn model described in Ref. Buchalla:2018yce , finding agreement between the two implementations at amplitude and cross section level. Example diagrams contributing at Born level are shown in Fig. 1.
Figure 1: Example diagrams contributing to Higgs plus one jet production at leading order, based on the chiral Lagrangian given in eq. (1).
@@
The Born matrix elements arising from the tree level diagrams where the Higgs boson couples to gluons, $\mathcal{M}^{\mathrm{0L}}_{Hg}$, are added to the rescaled SM Born matrix elements at form factor level, using the projectors of eqs. (9) and (10). The SM virtual two-loop amplitudes were constructed from about 2000 points computed numerically. They were then added at histogram level to the results obtained from the Born and real radiation contributions that are implemented in the POWHEG-BOX-V2.
@@
In order to allow for comparisons and cross checks, we implemented both the $m_{t}\to\infty$ limit as well as the full SM amplitudes at NLO. We checked that taking $m_{t}\to\infty$ in all diagrams and setting $c_{g}=0$ agrees with the SM calculation in the HTL. Furthermore, using the fact that in the HTL the SM reduces to diagrams with an effective Higgs-gluon coupling given by $c_{g,\mathrm{HTL}}=2/3$, see e.g. Spira:1995rr , and that the HEFT diagrams with a gluon-Higgs coupling reduce to just HTL diagrams without any top-loops, we checked that $c_{t}\mathcal{M}_{f}(m_{t}\to\infty)+c_{g}\mathcal{M}_{Hg}(m_{t}\to\infty)=(2/3c_{t}+c_{g})/c_{g,\mathrm{HTL}}\cdot\mathcal{M}_{\mathrm{HTL}}$. Of course, we also checked that taking $c_{t}=1$ and $c_{g}=0$ agrees with the SM results computed in Jones:2018hbb ; Chen:2021azt , both at amplitude level as well as at total cross section level.

### v2

@@ 1 Introduction
Run 2 of the LHC was very successful in establishing the Higgs boson couplings to vector bosons and heavy fermions. With Run 3 and the high-luminosity upgrade the precision of the coupling measurements will further increase. Therefore it is important to have good control over the theoretical uncertainties, in particular to distinguish beyond the Standard Model (BSM) effects from effects due to insufficient higher order corrections.
@@
It has been noticed some time ago Harlander:2013oja ; Azatov:2013xha ; Banfi:2013yoa ; Grojean:2013nya ; Schlaffer:2014osa ; Buschmann:2014twa ; Dawson:2014ora ; Buschmann:2014sia ; Langenegger:2015lra ; Azatov:2016xik ; Lindert:2018iug ; DiMicco:2019ngk that the Higgs boson $p_{T}$-spectrum is an important observable to constrain both the Yukawa-couplings to top (and bottom) quarks as well as effective gluon-Higgs couplings due to interactions with unknown particles at higher scales, leading to operators that are not present in the Standard Model (SM). The impact of such effective operators on the $p_{T,H}$ spectrum in Higgs+jet production, including next-to-leading Order (NLO) QCD corrections in the heavy top limit (HTL), has been investigated in Refs. Maltoni:2016yxb ; Grazzini:2016paz ; Grazzini:2018eyk ; Battaglia:2021nys ; Maltoni:2024dpn ; DiNoi:2024ajj , where the analysis of Refs. Battaglia:2021nys ; Maltoni:2024dpn ; DiNoi:2024ajj also includes investigations of effects due to the renormalisation group running of the Wilson coefficients.
@@
In the SM, the leading order (LO) one-loop amplitude has been calculated in Ref. Baur:1989cm , NLO results beyond the HTL were first obtained for the high transverse momentum region Melnikov:2016qoc ; Lindert:2018iug . The NLO result for general kinematics has first been obtained numerically Jones:2018hbb ; Becker:2020rjp ; Neumann:2018bsx , the top quark mass effects have been studied in detail in Ref. Chen:2021azt . Numerical results for the two-loop amplitude for Higgs plus jet production with full top quark mass dependence also have been calculated in Ref. Czakon:2021yub in the framework of the full next-to-next-to-leading order (NNLO) corrections to inclusive Higgs production, a study of the Higgs boson transverse momentum spectrum through the combination of this calculation with a parton shower has been presented in Ref. Niggetiedt:2024nmp .
@@
On the analytic side, compact analytic expressions for the one-loop amplitudes for Higgs + 4 parton scattering, providing an efficient way to calculate the NLO real radiation contributions, have been presented in Ref. Budge:2020oyl . After the relevant two-loop master integrals have become available Bonciani:2016qxi ; Bonciani:2019jyb ; Frellesvig:2019byn , the full NLO calculation based on analytic results for the two-loop integrals has been completed in Ref. Bonciani:2022jmb , including both top and bottom masses as well as a comparison of the on-shell and $\overline{\mathrm{MS}}$ schemes to renormalise the top quark mass. Two-loop bottom quark mass effects also have been studied in Ref. Pietrulewicz:2023dxt . Mixed QCD-EW corrections have been presented in Refs Bonetti:2020hqh ; Becchetti:2021axs ; Bonetti:2022lrk , partial EW corrections are considered in Refs. Becchetti:2018xsk ; Davies:2023npk ; Haisch:2024nzv . NNLO results for Higgs plus jet production in the HTL are available already since some time ago Boughezal:2015dra ; Boughezal:2015aha ; Caola:2015wna ; Chen:2016zka ; Campbell:2019gmd ; Chen:2019wxf ; Chen:2021ibm . Higgs production in association with one or multiple high-energy jets is available within the HEJ framework Andersen:2023kuj ; Andersen:2022zte . In Ref. Liu:2024tkc , light quark mediated Higgs boson production in association with a jet at NNLO and beyond is considered in the framework of resummation.
@@
The Higgs boson transverse momentum spectrum for boosted Higgs bosons already has been used by the experimental collaborations to place limits on anomalous top-Higgs and gluon-Higgs couplings ATLAS:2019lwq ; ATLAS:2021tbi ; CMS:2020zge ; ATLAS:2022fnp ; CMS:2024jbe . The latter can be parameterised by $c_{t}$ and $c_{g}$ in Higgs Effective Field Theory (HEFT), also called Electroweak Chiral Lagrangian Feruglio:1992wf ; Burgess:1999ha ; Grinstein:2007iv ; Contino:2010mh ; Alonso:2012px ; Buchalla:2013rka . These couplings also enter inclusive Higgs boson production in gluon fusion, which is known to agree with the SM prediction to a level approaching 5%. Therefore, these anomalous couplings are fairly well constrained already (also from other processes such as $t\mathrm{bar}{t}H$ production for the case of the Higgs-top coupling Hartland:2019bjb ; Celada:2024mcf ). However, it is well known that there is a degeneracy between $c_{t}$ and $c_{g}$ when considering only inclusive Higgs production, which is lifted when considering the $p_{T}$-spectrum of the Higgs boson at large transverse momenta Grojean:2013nya ; Schlaffer:2014osa ; Grazzini:2016paz ; Grazzini:2018eyk . Up to now, the effects of these operators have not yet been studied in combination with NLO corrections to Higgs+jet production including the full top quark mass dependence. However, both the SM top quark mass effects as well as these anomalous couplings affect the tail of the $p_{T,H}$-distribution considerably. Therefore it is important to study in detail the interplay of both, higher order QCD corrections and potential effects of new physics in an EFT framework.
@@
In this work we would like to address this point and investigate the effects of these anomalous couplings on the Higgs+jet cross section and transverse momentum distributions, based on a calculation of the full NLO QCD corrections in the SM Jones:2018hbb . Working in HEFT rather than Standard Model Effective Field Theory (SMEFT) Buchmuller:1985jz ; Grzadkowski:2010es ; Brivio:2017vri ; Isidori:2023pyp , our power counting scheme is not based on canonical dimension counting, i.e. the counting of inverse powers of a new physics scale $\Lambda$, but instead on the counting of the chiral dimension $d_{\chi}$, which is related to the (explicit or implicit) loop order $L$ through $d_{\chi}=2L+2$. The chromomagnetic top-quark dipole operator has chiral dimension $d_{\chi}=4$ and therefore is considered to be subleading, as explained in Section 2. Thus the inclusion of chromomagnetic dipole operators will be considered in subsequent work, together with other subleading operators entering at the same level, such as four-fermion operators.
@@
The structure of this paper is as follows: In Section 2, we describe how the anomalous couplings relate to inclusive Higgs production and how they affect the large-$p_{T}$ spectrum of the Higgs boson. We also give some technical details about the calculation. Section 3 is dedicated to the description of phenomenological results, providing heat maps that show the effects of the anomalous couplings on the total cross section and discussing the effect of some HEFT benchmark points on the Higgs boson transverse momentum spectrum, before we conclude in Section 4.
@@ 2.1 Framework of the calculation
In the SM, $c_{t}=1$ and $c_{g}=0$. We assume that the anomalous couplings are induced by new physics interactions at a scale $\Lambda$ considerably larger than the electroweak scale. While the process $pp\to H+$jet is loop induced in the SM, the second part of the Lagrangian now also introduces effective tree level interactions. The factor proportional to $\alpha_{s}/8\pi$ indicates that these interactions are stemming from loops of heavy particles that have been integrated out to arrive at the effective Higgs-gluon coupling. In the region $2m_{t}\lesssim\sqrt{\hat{s}}\lesssim\Lambda$ the top quark loops are resolved while the heavier particles in the loop generate the effective point-like Higgs-gluon interaction. The coefficient $c_{t}$ is a modification factor of the top Yukawa coupling, which can arise for example by mixing with heavy top partners Grojean:2013nya ; Banfi:2019xai . The chromomagnetic top-quark dipole operator ${\cal O}_{tg}=y_{t}g_{s}\mathrm{bar}t_{L}\sigma^{\mu\nu}T^{a}G_{\mu\nu}^{a}(v+H)t_{R}$ can only be generated through contracted loops Arzt:1994gp ; Buchalla:2022vjp ; Isidori:2023pyp in weakly coupled, renormalisable UV completions and we stick to such extensions of the SM. Therefore, inserting the operator into the one-loop diagrams that constitute the LO in the SM, the resulting diagrams are effectively of two-loop order (chiral dimension $d_{\chi}=6$), thus coming with a loop suppression factor $\sim 1/(16\pi^{2})$. The operator ${\cal O}_{hg}$, mediating direct Higgs-gluon couplings, also pertains to the class of loop-generated operators. However, insertions of ${\cal O}_{hg}$ enter at tree-level and therefore have chiral dimension $d_{\chi}=4$. The SM NLO QCD corrections also lead to two-loop diagrams and therefore come with a loop factor $1/(16\pi^{2})$ relative to the LO diagrams. However, the SM two-loop QCD diagrams come with an extra factor $g_{s}^{2}=4\pi\alpha_{s}$ relative to the Born diagrams, while the chromomagnetic dipole operator comes with an extra factor $y_{t}g_{s}$. As we only calculate the QCD corrections, which are of order ${\cal O}(\alpha_{s})$ relative to the Born amplitude, we neglect the two-loop corrections stemming from the chromomagnetic top-quark dipole operator in the present work. This is in line with the procedure followed in Ref. Buchalla:2018yce for the QCD corrections to Higgs boson pair production. Furthermore, we do not include any 4-quark operators nor CP-violating operators.
@@
where $\mathcal{M}_{f}$ denotes the parts of the amplitude where the Higgs boson couples to a top quark, and $\mathcal{M}_{Hg}$ the amplitude parts containing an effective point-like Higgs-gluon interaction. Note that the NLO amplitude can also contain both, a top quark loop and a point-like Higgs-gluon interaction. However, in these diagrams the top quarks couple only to gluons, with an SM coupling. We use $\kappa_{g}=\frac{3}{2}c_{g}$ such that the HTL corresponds to $\kappa_{g}\to 1$ and $c_{t}=0$. The total cross section for $pp\to H+\rm{jet}$ can be written as a quadratic polynomial in $c_{t}$ and $c_{g}\,$, both at LO and at NLO.
Integrating over the jet momenta, the total inclusive cross section for Higgs boson production in gluon fusion should be retrieved. As is well known, due to the “Higgs Low Energy Theorem” Ellis:1975ap ; Vainshtein:1980ea ; Dawson:1989yh , the total cross section for Higgs production in gluon fusion is rather insensitive to the masses of heavy particles circulating in the loop. This is also reflected in the fact that, at energy scales below $2m_{t}$, the inclusive Higgs production cross section is approximated very well by the HTL. An extra high-$p_{T}$ jet can serve as a handle to resolve heavy quark loops, therefore new physics effects could show up in the tail of the Higgs $p_{T}$-distribution.
@@
As the measured total Higgs production cross section in gluon fusion agrees very well with the SM result, the relation $(c_{t}+\kappa_{g})^{2}=1$ should be fulfilled to about 10% level, assuming that subleading operators do not have a drastic effect, which would lead to more freedom in the relation between $c_{t}$ and $\kappa_{g}$. In the pure HTL, the proportionality of the cross section to $(c_{t}+\kappa_{g})^{2}$ is fulfilled exactly, also for the NLO amplitudes, because there are no diagrams at NLO which contain both $c_{t}$ and $c_{g}$ simultaneously, such that the HTL of the full SM NLO amplitudes gives exactly the amplitudes proportional to $\kappa_{g}$. This degeneracy is broken in the Higgs boson transverse momentum spectrum because as $p_{T,H}$ increases, the top quark loops start to become resolved and therefore the kinematic behaviour of the contribution proportional to $c_{t}$ is different from the one in the HTL for large values of $p_{T,H}$. On the other hand, as the differential cross section decreases rapidly with $p_{T,H}$, the effects of anomalous couplings on the total cross section should be small as long as the relation $(c_{t}+\kappa_{g})^{2}=1$ is fulfilled. Therefore it is useful to consider the cross section for $pp\to H+\rm{jet}$ as a function of the cut on $p_{T,H}$ Grojean:2013nya ; Schlaffer:2014osa :
@@
where the coefficients $\delta_{i}$ depend on the cut $p_{T,H}^{\mathrm{min}}$. For small $p_{T,H}^{\mathrm{min}}$ the coefficients $\delta_{i}$ at LO are very small, modifying the cross section in the permille to percent range below $p_{T,H}\sim 350~{}\mathrm{G}\mathrm{eV}$ Grojean:2013nya . However, recent LHC measurements have reached transverse momentum regions beyond $600~{}\mathrm{G}\mathrm{eV}$ CMS:2024jbe ; CMS:2020zge ; ATLAS:2019lwq ; ATLAS:2023jdk . Furthermore, the study of Refs. Grojean:2013nya ; Schlaffer:2014osa was at LO only, and the one of Refs. Grazzini:2016paz ; Maltoni:2016yxb ; Grazzini:2018eyk is based on the HTL when going beyond LO. In Section 3, we will investigate how the anomalous couplings modify the large-$p_{T,H}$ spectrum at NLO with full top quark mass dependence.
@@ 2.2 Technical details
The LO amplitudes in the full theory as well as the amplitudes involving $c_{g}$ were implemented analytically, relying on Ref. Baur:1989cm , while the one-loop real radiation contribution and the two-loop virtual amplitudes rely on semi-numerical evaluations. As a cross-check we also generated the Born amplitudes with GoSam Cullen:2011ac ; GoSam:2014iqq using the UFO Degrande:2011ua ; Darme:2023jdn model described in Ref. Buchalla:2018yce , finding agreement between the two implementations at amplitude and cross section level. Example diagrams contributing at Born level are shown in Fig. 1.
Figure 1: Example diagrams contributing to Higgs plus one jet production at LO, based on the chiral Lagrangian given in eq. (1).
@@
The Born matrix elements arising from the tree level diagrams where the Higgs boson couples to gluons, $\mathcal{M}^{\mathrm{0L}}_{Hg}$, are added to the rescaled SM Born matrix elements at form factor level, using the projectors of eqs. (9) and (10).
In order to use the 2-loop virtual contribution directly within the POWHEG-BOX-V2, we constructed a grid based on the 2000 phase space points at which the 2-loop amplitude has been evaluated, together with an interpolation framework. To this aim we first realise that we can write the amplitude as
$M(\mu)=M_{2}\log(\mu)^{2}+M_{1}\log(\mu)+M_{0}.$ (12)
Thus, we computed the amplitudes for three values of $\mu$ and then solved the equations for $M_{0},M_{1}$ and $M_{2}$. We created a grid for each of the four partonic channels where we stored
$\beta=\frac{s-m_{H}^{2}}{s+m_{H}^{2}},~{}\cos\theta=\frac{t-u}{s-m_{H}^{2}},~{}M_{0},M_{1},M_{2}.$ (13)
To interpolate the virtual amplitudes for each channel we used a neural network trained on the grid, where we used 30% of the data points as a validation set. For the training of the network we multiplied the coefficients $M_{0},M_{1}$ and $M_{2}$ with $\beta^{2}(1-\beta)(1-\cos^{2}\theta)$ and divided by the largest value in the grid. This guarantees that the grid for the neural network training has values in $[-1,1]$ and is flattened at the phase space boundaries. The architecture of each network is composed of three dense layers with 200, 20 and 20 nodes, respectively. The setup is implemented using Keras chollet2015keras to produce the models in combination with a modified version of Keras2cpp keras2cpp to save them such that they can be loaded from C++. We built a C++ function where, for each channel, the models are loaded. This function is called in the POWHEG-BOX-V2 to obtain the two-loop virtual amplitude at any given phase space point.
@@
One full run takes roughly 320 CPU-days, depending on the used hardware. The 320 CPU-days refer to a cluster which consists of Intel Xeon Gold 6230 processors with a frequency of 2.1 GHz. It is also worth mentioning that using either $c_{t}=0$ or $c_{g}=0$ can drastically reduce the number of CPU-hours, since this greatly reduces the number of diagrams to be evaluated.
@@
In order to allow for comparisons and cross checks, we implemented both the $m_{t}\to\infty$ limit as well as the full SM amplitudes at NLO. We checked that taking $m_{t}\to\infty$ in all diagrams and setting $c_{g}=0$ agrees with the SM calculation in the HTL. Furthermore, using the fact that in the HTL the SM reduces to diagrams with an effective Higgs-gluon coupling given by $c_{g,\mathrm{HTL}}=2/3$, see e.g. Spira:1995rr , and that the HEFT diagrams with a gluon-Higgs coupling reduce to just HTL diagrams without any top-loops, we checked that $c_{t}\mathcal{M}_{f}(m_{t}\to\infty)+c_{g}\mathcal{M}_{Hg}(m_{t}\to\infty)=(2/3c_{t}+c_{g})/c_{g,\mathrm{HTL}}\cdot\mathcal{M}_{\mathrm{HTL}}$. Furthermore, we have validated the interpolation grid that we implemented in the Powheg setup by comparing the grid-based result with the result reconstructed directly from the points obtained from the numerical evaluation of the two-loop amplitude. The differences were of the order of the Monte Carlo uncertainties. We also interpolated the Born amplitude of the grid and compared it to the analytical results. Of course, we also checked that taking $c_{t}=1$ and $c_{g}=0$ agrees with the SM results computed in Jones:2018hbb ; Chen:2021azt , both at amplitude level and at total cross section level.